Watermate Products is a manufacturer of large flower pots for urban settings. The company has these​ standards:

Direct materials (resin) - 10.8 pounds per pot, at a cost of $3.00 per pound

Direct labor - 1.6 hours at a cost of $15.65 per hour

Standard variable manufacturing overhead rate - $2.50 per direct labor hour

Predetermined fixed manufacturing overhead rate - $6.60 per direct labor hour

1.

Compute the standard cost of each of the following inputs per​ pot: direct​ materials, direct​ labor, variable manufacturing​ overhead, and fixed manufacturing overhead.

2.

Determine the standard cost of one flower pot.
Answer #1

1. Standard cost:

Direct materials (resin) = 10.8*3 = $32.4

Direct labor = 1.6*15.65 = $25.04

variable manufacturing overhead rate = 1.6*2.50 = $4.00

Fixed manufacturing overhead = 1.6*6.60 = $10.56

2. Standard cost of one flower pot = 32.4+25.04+4+10.56 = $72